### v2

@@ 8 Query lower bound on adaptively testing sparsity
Let $\mathbb{I}_{x}$ be the indicator variable that takes the value 11 when $f(x)\neq g(x)$, and it is 0 otherwise. Then $X_{g}^{\prime}=\frac{\sum_{x\in\mathbb{Z}_{p_{i_{1}}^{m_{i_{1}}}}\times\cdots\times\mathbb{Z}_{p_{i_{Ct}}^{m_{i_{Ct}}}}}\mathbb{I}_{x}}{p_{i_{1}}^{m_{i_{1}}}\cdots p_{i_{Ct}}^{m_{i_{Ct}}}}$ denotes the Hamming distance between them. Also, by Lemma 2.43, if $X_{g}$ denotes thesquare of the $\ell_{2}$ distance between $f$ and $g$, then $X_{g}=4X_{g}^{\prime}$. Now,
@@
Let AA be the set of all functions $f:\mathbb{Z}_{p_{i_{1}}^{n_{i_{1}}}}\times\cdots\times\mathbb{Z}_{p_{i_{Ct}}^{n_{i_{Ct}}}}\to\{-1,+1\}$ which are $(2-\tau)$ close to $\deg_{\mathcal{G}}=t$ in $\ell_{2}$. Then $\mathcal{P}_{No}(A)=0$, and by Lemma 8.6, $\mathcal{P}_{Uniform(A)}\leq\frac{1}{10}$, where $\mathcal{P}_{No}$ and $\mathcal{P}_{Uniform}$ denote the probability over the distributions $\mathcal{D}_{No}$ and $\mathcal{D}_{Uniform}$ respectively. So, $|\mathcal{L}_{No}(A)-\mathcal{L}_{Uniform}(A)|\leq\frac{1}{10}$.

### v3

@@ 8 Query lower bound on adaptively testing sparsity
Let $\mathbb{I}_{x}$ be the indicator variable that takes the value $1$ when $f(x)\neq g(x)$, and it is 0 otherwise. Then $X_{g}^{\prime}=\frac{\sum_{x\in\mathbb{Z}_{p_{i_{1}}^{m_{i_{1}}}}\times\cdots\times\mathbb{Z}_{p_{i_{Ct}}^{m_{i_{Ct}}}}}\mathbb{I}_{x}}{p_{i_{1}}^{m_{i_{1}}}\cdots p_{i_{Ct}}^{m_{i_{Ct}}}}$ denotes the Hamming distance between them. Also, by Lemma 2.43, if $X_{g}$ denotes thesquare of the $\ell_{2}$ distance between $f$ and $g$, then $X_{g}=4X_{g}^{\prime}$. Now,
@@
Let $A$ be the set of all functions $f:\mathbb{Z}_{p_{i_{1}}^{n_{i_{1}}}}\times\cdots\times\mathbb{Z}_{p_{i_{Ct}}^{n_{i_{Ct}}}}\to\{-1,+1\}$ which are $(2-\tau)$ close to $\deg_{\mathcal{G}}=t$ in $\ell_{2}$. Then $\mathcal{P}_{No}(A)=0$, and by Lemma 8.6, $\mathcal{P}_{Uniform(A)}\leq\frac{1}{10}$, where $\mathcal{P}_{No}$ and $\mathcal{P}_{Uniform}$ denote the probability over the distributions $\mathcal{D}_{No}$ and $\mathcal{D}_{Uniform}$ respectively. So, $|\mathcal{L}_{No}(A)-\mathcal{L}_{Uniform}(A)|\leq\frac{1}{10}$.